OpenStudy (anonymous):

it is given that y varies directly as (ax-4) where a is a non zero constant.y=16 when x=4 and y=34 when x=7, find the value of a

OpenStudy (bahrom7893):

y = k(ax-4) y = 16 when x=4: 16=k(a*4-4)

OpenStudy (bahrom7893):

y=34 when x=7: 34=k(a*7-4)

OpenStudy (bahrom7893):

Solve the system of equations: 16=k(4a-4) 34=k(7a-4) That will get you your answer.

OpenStudy (anonymous):

thank you~